# bibliography/references.bib *[itemize]labelindent=10pt, itemindent=0pt, leftmargin=*

### v1

@@ Abstract.
High-Energy Physics experiments are facing a multi-fold data increase with every new iteration. This is certainly the case for the upcoming High-Luminosity LHC upgrade. Such increased data processing requirements forces revisions to almost every step of the data processing pipeline. One such step in need of an overhaul is the task of particle track reconstruction, a.k.a., tracking. A Machine Learning-assisted solution is expected to provide significant improvements, since the most time-consuming step in tracking is the assignment of hits to particles or track candidates. This is the topic of this paper.
@@ 1. Introduction
Tracking refers to the reconstruction of the trajectory (track) of subatomic particles present in particle physics experiments. The task of tracking involves reconstructing the trajectory of a particle based on sensory data. A sensor may or may not be hit by a particle, and sequences of these hits comprise tracks. In HEP, the sensory data can come from detectors such as, ALICE (Collaboration:2008:ALICE), ATLAS (Collaboration:2008:ATLAS), CMS (Collaboration:2008:CMS), and LHCb (Collaboration:2008:LHCb), installed at the Large Hadron Collider (LHC). At the same time, tracking is important for experiments in neutrino physics, astroparticle physics and other fields, e.g., (Woodruff:2018:APTI; Aguilar:2011:FAMT; Patterson:2009:ETER; González-Díaz:2015:AGMT).
@@ 1.1. Our contribution
This paper focuses on the development of first solutions for the assignment of hits to track/particle candidates using the Transformer and the U-Net ML model architectures. The challenge is to achieve high solution accuracy, which corresponds to the accuracy of assigning hits to correct tracks, while achieving better computational efficiency. Here, computational efficiency is measured by the mean execution time per event.
@@
Our goal is to offer fast, i.e., computationally-efficient and high-throughput, solutions for tracking. Therefore, we strive for models that assign hits to trajectories in a single inference step. We determine the accuracy and speed of these models for the tracking problem and use simulated data of increasing complexity to compare our approaches. Finally, we discuss how our initial approaches can improve their performance through post-processing steps and argue that our approaches are fast and accurate enough to be considered as candidates for future tracking pipelines.
@@ 1.2.2. Simulations for HEP
As HEP experiments are not of the kind to be performed on demand, simulations are the next best thing. Simulations for HEP can be used to study the effects of physics phenomena through the generation of data sets for analyses and algorithm design efforts. There are numerous simulations available, predominantly focusing on physics-accuracy and detector specificity. Examples relevant to the ATLAS detector are Geant4 (Agostinelli:2003:GST), FATRAS (Edmonds:2008:FATS) and ATLFAST (Richter-Was:1998:AFSP).
@@ 1.2.3. Traditional methods and performance
Variations of Kalman filtering have been used at LHC for decades (Kalman:1960:NALF; Frühwirth:1987:AKFT; Braun:2019:CKFH) due to its robustness and excellent performance. However, its combinatorial nature in the track finding stage and inherent sequential execution make KF unsuitable for the upcoming High-Luminosity stage of the LHC, as it scales poorly with the number of recorded hits. The currently utilised algorithm is tested on simulated data of events with 200 points of origin of tracks (pile-up) (Atlas:2019:FTRH), which is similar in complexity to the largest data evaluated in this paper. The KF pipeline takes 214.3 HS06 × seconds for a single event, translating to around 12 seconds CPU-time [^1] The CPU-time is multiplied by the HS06 factor of 17.8 for single-threaded execution, since the considered CPU is an Intel Xeon E5-2620v2. HS06 is a benchmark for measuring CPU performance in HEP. More information can be found at: https://w3.hepix.org/benchmarking/HS06.html. An optimised algorithm with tighter track selection in the track finding stage achieves a 7x speed-up, requiring 1.8 seconds of CPU-time per event (Atlas:2019:FTRH).
@@ 1.2.4. Machine learning competitions and community engagement
To engage data scientists and machine learning experts, a tracking competition (TrackML) (Kiehn:2019:TrackML) was launched on the Kaggle platform. A training data set was created based on the simulation of a generic High-Luminosity LHC (HL-LHC) experiment tracker, listing the measured 3D hits/points for each event and the list of 3D hits belonging to a real track. The events are top-pair events with a ”pile-up” of $\mu=200$ using Poisson statistics for the superposition of QCD events.
@@ 1.2.5. Advances in deep learning and GNNs
Given that particle interactions can be conveniently represented as graphs, Graph Neural Networks (GNNs) have emerged as a promising approach. These networks operate on graph-structured data, capturing the relational information between hits in a detector. The state-of-the-art in recent years has focused on GNN-based solutions (Bakina:2022:DLTR; Biscarat:2021:TRTR; Choma:2020:TSLE; DeZoort:2021:CPTE; Duarte:2022:GNNP; Elabd:2022:GNNC; Farrell:2018:NDLM; Heintz:2020:ACPT; Ju:2020:GNNP; Ju:2021:PGDL; Mieskolainen:2023:HNCH; Murnane:2023:EGNN; Thais:2021:ISGO; Duarte:2022:GNNP), or this reference for a review on GNNs in HEP (Shlomi:2021:GNNP).
In GNN-based methods, edges between vertices (hits) are predicted to determine actual physical trajectories. A graph is generated based on an event, with all hits as nodes connected by edges based on some constraint [^2] For example, geometric constraints or a preprocessing algorithm such as Hough Transform (Tsaris:2018:TRKX).. A GNN is trained to assign weights to edges or prune unlikely ones, ensuring a fitting connectivity between vertices to represent particle trajectories (DeZoort:2021:CPTE; Ju:2021:PGDL; Lazar:2023:AIEP).
A recent successful approach (Lieret:2024:OCPC; Lieret:2023:HPPT) involves constructing a graph by connecting hits from different detector layers that satisfy certain geometric constraints. A fully connected neural network estimates the weights of all edges, pruning those below a threshold. Next, object condensation clusters hits of the same track in a learned space, regressing the properties of the reconstructed objects (Kieseler:2020:OCOG). Clustering is done using Density-Based Spatial Clustering of Applications with Noise (DBSCAN). Performance is assessed on the inner detector hits from the TrackML data. Lieret et al. (Lieret:2023:HPPT) defined multiple custom metrics to assess physics performance. No time performance is reported, but the authors suggest that using Transformer models would significantly reduce inference time. An interesting benchmark for GNNs is (Ju:2021:PGDL), which reports an inference time of 2.2 seconds wall-clock time (including data transfer to GPUs) on the full TrackML events, using an NVIDIA A100 GPU. This approach involved significant pre and post-processing of the data and reports a track ML accuracy of about $0.87$.
Finally, we would like to point out that transformers have already been used in many applications in HEP and have proven their versatility and efficiency (Qu:2022:PTJT; Mikuni:2021:PCTC; Butter:2023:JDMN; Leigh:2024:DPCG; Builtjes:2024:ASPI; Caron:2024:UADL).
@@ 1.2.7. Data reduction and standardisation
The most notable trend is the de facto use of TrackML (Amrouche:2021:TMLC) data for experimentation. However, ML model training is computationally expensive and requires large amounts of hardware resources, such as GPU memory. Therefore, authors often reduce the data. The most common reduction is to only consider data associated with the inner detector, also known as the pixel detector (Choma:2020:TSLE; DeZoort:2021:CPTE; Elabd:2022:GNNC; Heintz:2020:ACPT; Ju:2020:GNNP; Murnane:2023:EGNN; Thais:2021:ISGO). A few authors apply further reductions, such as noise hit reduction (Choma:2020:TSLE; Ju:2020:GNNP) or filtering for limited $P_{T}$ values (Heintz:2020:ACPT). Some examples consider the full detector but apply lower pile-up alongside noise hit reduction (Mieskolainen:2023:HNCH).
@@ 2. Data sets
In this section, we describe the data sets used to train and evaluate our particle tracking methods. These data sets cover a spectrum of scale and track representation complexity, ranging from simple linear tracks to more complex helical and real-world tracks. By increasing problem complexity in two dimensions—scale and track representation—we can thoroughly assess the performance and robustness of the tracking algorithms. We consider five data sets, covering a spectrum of scale and track representation complexity from low to high,
@@
  • 10-50 (variable count) tracks per event from the TrackML data set,
@@
  • 200-500 (variable count) tracks per event from the TrackML data set.
@@
The first three data sets are the result of simulations using REDVID simulation framework (Odyurt:2024:RSHE). REDVID simulations are fully configurable with an extensive set of options. We have considered events with random track counts. Simulations with minimum and maximum track count boundaries as $[10,50]$ and $[50,100]$ have been executed. Track function complexity varies between linear and expanding helical, with the latter representing a simple emulation of charged particles in a magnetic field. Note that the 3D geometric space and elements contained within are defined in cylindrical coordinate system, with $r$, $\theta$ and $z$ coordinates as radius, angle with the X-axis and location on the Z-axis, respectively.
@@
To move forward in the complexity spectrum, beyond the “50-100 helical tracks REDVID” data set, we have opted to switch to the data associated with the TrackML Kaggle challenge (Kiehn:2019:TrackML). Reduced versions of this data set are frequently used to assess the state-of-the-art models for this task. Released in 2018, the goal of this challenge was to identify machine learning solutions to the track reconstruction problem. For that purpose, a huge data set representing the conditions in the High-Luminosity stage of the LHC was created.
@@
The TrackML data set consists of 3D hit coordinates in a Cartesian coordinate system, with the global Z-axis defined along the beam direction (Amrouche:2021:TMLC). There are ten available track parameters, but we only make use of four of them as necessary to uniquely identify tracks: the particle charge $q$, and the initial momentum (in GeV/c) along each global axis ($p_{x},p_{y},p_{z}$). However, the momenta are transformed into the spherical coordinate system according to
@@
TrackML statistics
Though resulting from a simulation, the TrackML data has considerable scale. It contains 8 850 individual events with an average of 10 780 particles, between 6 898 and 16 197, per event, which could be interpreted as the average track count. It includes particles represented with as many as 28 hit points. At the same time, there are particles with 0 associated hit points. The average particle hit count throughout the events is 8.52. Event collision density varies from 1300 to 3048. Substantial amount of noise hits are present, which is one of the main factors making TrackML a challenging data set. Almost every publication considering this data set has performed some sort of noise reduction. For our highest complexity data set, “200-500 tracks per event from the TrackML data set”, we have selected between 200 to 500 tracks at random per event. This process is repeated five times, resulting in 43 725 reduced events.

### v2

@@ Abstract.
High-Energy Physics experiments are facing a multi-fold data increase with every new iteration. This is certainly the case for the upcoming High-Luminosity LHC upgrade. Such increased data processing requirements force revisions to almost every step of the data processing pipeline. One such step in need of an overhaul is the task of particle track reconstruction, a.k.a., tracking. A Machine Learning-assisted solution is expected to provide significant improvements, since the most time-consuming step in tracking is the assignment of hits to particles or track candidates. This is the topic of this paper.
@@ 1. Introduction
Tracking refers to the reconstruction of the trajectory (track) of subatomic particles present in particle physics experiments. The task of tracking involves reconstructing the trajectory of a particle based on detector data. A sensor may or may not be hit by a particle, and sequences of these hits comprise tracks. In HEP, the sensory data can come from detectors such as, ALICE (Collaboration:2008:ALICE), ATLAS (Collaboration:2008:ATLAS), CMS (Collaboration:2008:CMS), and LHCb (Collaboration:2008:LHCb), installed at the Large Hadron Collider (LHC). At the same time, tracking is important for experiments in neutrino physics, astroparticle physics and other fields, e.g., (Woodruff:2018:APTI; Aguilar:2011:FAMT; Patterson:2009:ETER; González-Díaz:2015:AGMT).
@@ 1.1. Our contribution
This paper focuses on the development of first solutions for the assignment of hits to track/particle candidates using the Transformer and the U-Net ML model architectures. Note that for our current solutions, we have opted to address the formation of hit clusters matching tracks and do not intend to cover the actual parameterisation of a track function, which would be a separate non-trivial task. The challenge is to achieve high solution accuracy, which corresponds to the accuracy of assigning hits to correct tracks, while achieving better computational efficiency. Here, computational efficiency is measured by the mean execution time per event.
@@
Our goal is to offer fast, i.e., computationally-efficient and high-throughput, solutions for tracking. Therefore, we strive for models that assign hits to trajectories in a single inference step. We determine the accuracy and speed of these models for the tracking problem and use simulated data of increasing complexity to compare our approaches. Finally, we discuss how our initial approaches can improve their performance through post-processing steps and argue that our approaches are fast and accurate enough to be considered as candidates for future tracking pipelines. Our data sets (Odyurt:2024:DATASET) and code (Dobreva:2024:CODE) [^1] Also available at https://github.com/TrackFormers/TrackFormers-Models are publicly available.
@@ 1.2.2. Simulations for HEP
As HEP experiments are not of the kind to be performed on demand, simulations are suitable and often necessary alternative. Simulations for HEP can be used to study the effects of physics phenomena through the generation of data sets for analyses and algorithm design efforts. There are numerous simulations available, predominantly focusing on physics-accuracy and detector specificity. Examples relevant to the ATLAS detector are Geant4 (Agostinelli:2003:GST), FATRAS (Edmonds:2008:FATS) and ATLFAST (Richter-Was:1998:AFSP).
@@ 1.2.3. Traditional methods and performance
Variations of Kalman filtering have been used at LHC for decades (Kalman:1960:NALF; Frühwirth:1987:AKFT; Braun:2019:CKFH) (and similar GPU-based algorithms (Aaij:2020:ALLEN)) due to its robustness and excellent performance. However, its combinatorial nature in the track finding stage and inherent sequential execution make KF unsuitable for the upcoming High-Luminosity stage of the LHC, as it scales poorly with the number of recorded hits. The currently utilised algorithm is tested on simulated data of events with 200 points of origin of tracks (pile-up) (Atlas:2019:FTRH), which is similar in complexity to the largest data evaluated in this paper. The KF pipeline takes 214.3 HS06 × seconds for a single event, translating to around 12 seconds CPU-time [^2] The CPU-time is multiplied by the HS06 factor of 17.8 for single-threaded execution, since the considered CPU is an Intel Xeon E5-2620v2. HS06 is a benchmark for measuring CPU performance in HEP. More information can be found at: https://w3.hepix.org/benchmarking/HS06.html. An optimised algorithm with tighter track selection in the track finding stage achieves a 7x speed-up, requiring 1.8 seconds of CPU-time per event (Atlas:2019:FTRH).
@@ 1.2.4. Machine learning competitions and community engagement
To engage data scientists and machine learning experts, a tracking competition (TrackML) (Kiehn:2019:TrackML) was launched on the Kaggle platform. A training data set was created based on the simulation of a generic High-Luminosity LHC (HL-LHC) experiment tracker, listing the measured 3D hits/points for each event and the list of 3D hits belonging to a real track. The events are top-pair events with a ”pile-up” of $\mu=200$ using Poisson statistics for the superposition of Quantum ChromoDynamics (QCD) events, in which, there are concentrations of events in close proximity.
@@ 1.2.5. Advances in deep learning and GNNs
Given that particle interactions can be conveniently represented as graphs, Graph Neural Networks (GNNs) have emerged as a promising approach. These networks operate on graph-structured data, capturing the relational information between hits in a detector. The state-of-the-art in recent years has focused on GNN-based solutions (Bakina:2022:DLTR; Biscarat:2021:TRTR; Choma:2020:TSLE; DeZoort:2021:CPTE; Duarte:2022:GNNP; Elabd:2022:GNNC; Farrell:2018:NDLM; Heintz:2020:ACPT; Ju:2020:GNNP; Ju:2021:PGDL; Mieskolainen:2023:HNCH; Murnane:2023:EGNN; Thais:2021:ISGO; Duarte:2022:GNNP; Caillou:2024:PPAG), or see this reference for a review on GNNs in HEP (Shlomi:2021:GNNP).
In GNN-based methods, edges between vertices (hits) are predicted to determine actual physical trajectories. A graph is generated based on an event, with all hits as nodes connected by edges based on some constraint [^3] For example, geometric constraints or a preprocessing algorithm such as Hough Transform (Duda:1972:HTLC).. A GNN is trained to assign weights to edges or prune unlikely ones, ensuring a fitting connectivity between vertices to represent particle trajectories (DeZoort:2021:CPTE; Ju:2021:PGDL; Lazar:2023:AIEP).
A recent successful approach (Lieret:2024:OCPC; Lieret:2023:HPPT) involves constructing a graph by connecting hits from different detector layers that satisfy certain geometric constraints. A fully connected neural network estimates the weights of all edges, pruning those below a threshold. Next, object condensation clusters hits of the same track in a learned space, regressing the properties of the reconstructed objects (Kieseler:2020:OCOG). Clustering is done using Density-Based Spatial Clustering of Applications with Noise (DBSCAN). Performance is assessed on the inner detector hits from the TrackML data. Lieret et al. (Lieret:2023:HPPT) defined multiple custom metrics to assess physics performance. No time performance is reported, but the authors suggest that using Transformer models would significantly reduce inference time. An interesting benchmark for GNNs is (Ju:2021:PGDL), which reports an inference time of 2.2 seconds wall-clock time (including data transfer to GPUs) on the full TrackML events, using an NVIDIA A100 GPU. This approach involved significant pre and post-processing of the data and reports a TrackML score of about $0.87$. In our approach, we consider an iteration of the TrackML score, i.e., FitAccuracy, as a metric for scoring the success rates of our solutions. FitAccuracy score and its relation with TrackML score is described in detail in Section 4.
Finally, we would like to point out that transformers have already been used in many applications in HEP and have proven their versatility and efficiency (Qu:2022:PTJT; Mikuni:2021:PCTC; Butter:2023:JDMN; Leigh:2024:DPCG; Builtjes:2024:ASPI; Caron:2024:UADL). Though there are more visionary efforts towards the use of Transformers for foundational models, or to further generalise their application in collider physics, these efforts are not as mature yet (Huang:2024:LMPT).
@@ 1.2.7. Data reduction and standardisation
The most notable trend is the de facto use of TrackML (Amrouche:2020:TMLC) data for experimentation. However, ML model training is computationally expensive and requires large amounts of hardware resources, such as GPU memory. Therefore, authors often reduce the data. The most common reduction is to only consider data associated with the inner detector, also known as the pixel detector (Choma:2020:TSLE; DeZoort:2021:CPTE; Elabd:2022:GNNC; Heintz:2020:ACPT; Ju:2020:GNNP; Murnane:2023:EGNN; Thais:2021:ISGO). A few authors apply further reductions, such as noise hit reduction (Choma:2020:TSLE; Ju:2020:GNNP) or filtering for limited $P_{T}$ values (Heintz:2020:ACPT). Some examples consider the full detector but apply lower pile-up alongside noise hit reduction (Mieskolainen:2023:HNCH).
@@ 2. Data sets
In this section, we describe the data sets used to train and evaluate our particle tracking methods. These data sets cover a spectrum of scale and track representation complexity, ranging from simple linear tracks to more complex helical and closer to real-world tracks. By increasing the problem complexity in two dimensions (scale and track representation) we can thoroughly assess the performance and robustness of the tracking algorithms. As such, we consider five data sets, covering a spectrum of scale and track representation complexity from low to high. Data set titles are,
@@
  • 10-50 (variable count) tracks per event, extracted from the TrackML data set,
@@
  • 200-500 (variable count) tracks per event, extracted from the TrackML data set.
2.1. REDVID data set
@@
The first three data sets are the result of simulations using REDVID simulation framework (Odyurt:2024:RSHE). REDVID simulations are fully configurable with an extensive set of options. We have considered events with random track counts. Simulations with minimum and maximum track count boundaries as $[10,50]$ and $[50,100]$ have been executed. Track function complexity varies between linear and expanding helical, with the latter representing a simple emulation of charged particles in a magnetic field. Note that the 3D geometric space and elements contained within are defined in cylindrical coordinate system, with $r$, $\theta$ and $z$ coordinates as radius, angle with the X-axis and location on the Z-axis, respectively. The above REDVID data sets contain 100 000 events each, with noise (smearing) applied to calculated hit coordinates. Data set headers include 15 fields, covering
  • event, sub-detector, track and hit ids,
  • track function parameters, whether linear or helical, same number of parameters are present, with different interpretations per track type,
  • hit coordinates, and
  • a few type descriptors for sub-detector and track types.
For more detailed descriptions covering field data types, we refer readers to the README file in the shared Zenodo record (Odyurt:2024:DATASET)
2.1.1. REDVID detector geometry
At its core, a detector model is comprised of the geometric definitions of the included elements, shapes, sizes, and placements in space. Although we can support a variety of detector geometries, the overall structure, especially for our experimental results, resembles the ATLAS detector. Accordingly, there are four sub-detector types, Pixel, Short-strip, Long-strip and Barrel. The pixel and the barrel types have cylindrical shapes with the pixel being a filled cylinder, while the barrel being a cylinder shell with open caps. These are not hard requirements, as the geometry is fully parametric, and differing definitions can be opted for, e.g., a pixel as a cylinder shell. The long-strip and the short-strip types are primarily intended as flat disks, but can be defined as having a thickness, rendering them as cylinders. Sub-detector types can be selectively present or absent. Figure 1 depicts a representative variation of the detector geometry involving the aforementioned elements.
Figure 1. The fully parametric detector geometry, allowing for inclusion/exclusion of different sub-detector types, with full control over sub-layer counts, sizes and placements.
Structurally speaking, in a real-world detector, like ATLAS, the internals of short-strip and long-strip sub-detector types are different. We on the other hand, reduce such complexities to placement location and size, i.e., distance from the origin and sub-detector disk radius. Note that our geometric model does support disk thickness, which basically would turn disks into shallow cylinders. However, we have considered flat disks for our experiments.
2.1.2. REDVID event simulation
One of the simplifications for our complexity reduction approach is to consider a single collision per event (no pile-up), either aligning exactly to the origin point of the detector geometry, or at the presence of smearing for this alignment. This can be configured. However, the list of complexities, even without the polluting effects of multiple collisions, is extensive. Particles travelling through the detector matter could lead to secondary collisions, resulting in drastic changes in their trajectory. Such secondary collisions could also lead to the release of particles not originating from the collision event itself. These will show up as tracks with unusual starting points within the detector space, rather distant from the collision point. Some particles could also come to a halt, which would be seen as abruptly terminating tracks. Such physics-accurate behaviour of particles interacting with the present matter in detectors is not considered for our simulator. It must be noted that the generation of tracks originating far away from the origin and prematurely terminating tracks, can be emulated in our simulator in a randomised fashion.
2.2. TrackML data set
@@
To move forward in the complexity spectrum, beyond the “50-100 helical tracks REDVID” data set, we have opted to switch to the data associated with the TrackML Kaggle challenge (Kiehn:2019:TrackML). Reduced versions of this data set are frequently used to assess the state-of-the-art models for this task. Released in 2018, the goal of this challenge was to identify machine learning solutions to the track reconstruction problem.
The TrackML data set is a simulated collection of proton-proton collision events. It simulates specifically the production of top quark-antiquark pairs in the high-pileup conditions expected at the HL-LHC, with each event incorporating an average of 200 pileup interactions. The simulated detector is designed as a generalized representation of an LHC tracker, featuring discrete layers of cylinder and disk-shaped sensor arrays operating within a magnetic field as layers. The data set provides 3D hit positions along with truth-level information about the particles responsible for generating these hits. Some simplifications with respect to more realistic data were made. For example, it does not account for reconstructed hit merging, as such occurrences are rare (less than 0.5%). Moreover, it excludes complex smaller detector elements such as electronic systems, cooling tubes, or cables.
@@
The TrackML data set consists of 3D hit coordinates in a Cartesian coordinate system, with the global Z-axis defined along the beam direction (Amrouche:2020:TMLC). The data set contains ten values associated with every particle: particle identifier $pid$, the vertex of origin ($v_{x},x_{y},x_{z}$), the initial momentum (in GeV/c) along each global axis ($p_{x},p_{y},p_{z}$), the particle charge $q$ and the number of hits this particle generated $nhits$. We make use of only four of those as track defining parameters: $q,p_{x},p_{y},p_{z}$. However, the momenta are transformed into the spherical coordinate system according to
@@
Though resulting from a simulation, the TrackML data has considerable scale. It contains 8 850 individual events with an average of 10 780 particles, between 6 898 and 16 197, per event, which could be interpreted as the average track count. It includes particles represented with as many as 28 hit points. At the same time, there are particles with 0 associated hit points. The average particle hit count throughout the events is 8.52. The number of individual particle vertices composing a bunch crossing varies from 1300 to 3048. Substantial amount of noise hits are present (between 18 590 and 20 518 per event), which is one of the main factors making TrackML a challenging data set. Almost every publication considering this data set has performed some sort of noise reduction. For our highest complexity data set, “200-500 tracks per event from the TrackML data set”, we have selected between 200 to 500 tracks at random per event. This process is repeated five times, resulting in 43 725 reduced events.
The TrackML data set is described extensively in literature, for further information please refer to (Kiehn:2019:TrackML) or (Amrouche:2020:TMLC).
@@ 3.1. Transformer designs
3.2. U-Net design
The U-Net (Ronneberger:2015:UNet) is a CNN architecture that primarily segments images. It aims to classify each pixel in an image and allows it to distinguish and separate objects in complicated visual data. The ”U” in U-Net represents its U-shaped structure: it has a contracting path (encoder) and a wide-ranging path (decoder) with a bottleneck in between. On the one hand, contextual information is caught by the contracting path through the progressive downsampling of the input image using convolutional layers and pooling operations that reduce the spatial dimensions and increase feature abstraction. This process allows the network to grasp the global structure of objects in the image while improving its analytical capabilities. The most abstracted features are extracted and the smallest resolution is reached at the bottleneck. On the other hand, the decoder part aims to reconstruct the image at the original resolution by employing transposed convolutions or upsampling layers. It gradually refines the output by combining contextual information from the encoder with finer details provided by the skip connections thereby ensuring that the spatial details lost during downsampling are satisfactorily recovered.
Although U-Net architectures are being widely used for image segmentation tasks it is important to note that traditional convolutional networks usually work on densely populated spaces. For most of the ordinary tasks this does not pose a real problem in terms of memory usage and computational effort. However, in problems where the nature of the data involves relatively sparse information spanning along very large physical spaces, modifications to traditional CNN may be considered. These adaptations can help mitigate physical memory limitations in hardware and accelerate the mathematical operations performed. More recently, a novel convolution operator named as sparse convolution (SC) has been introduced in order to work with sparse data which may represent a suitable option for processing event detection data in three-dimensional geometries, such as those proposed in this manuscript. Interested readers about how sparse convolutions work and how they are implemented are referred to (Graham:2017:SSCN) and the references therein.